OpenStudy (anonymous):

medal will be awarded Find the sum of the summation of 3 i plus 2, from i equals 5 to 13

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

answer choices 261 41 220 38

OpenStudy (kropot72):

This question can be treated a an arithmetic series, with the first term being 15 + 2, the last term being 39 + 2, and the number of terms being 13 - 5 = 9. The required formula is \[\large S=\frac{n}{2}(a+L)\] where S is the sum of n terms, a is the first term and L is the last term.

OpenStudy (anonymous):

the first term is 13 and last term is 5 and the sum of terms 3i+2

OpenStudy (kropot72):

No. The first term is found by plugging in the value i = 5 into 3i + 2, giving: First term = (3 * 5) + 2 = 17 The last term is found by plugging in the value i = 13 into 3i + 2, giving: Last term = (3 * 13) + 2 = 41 Do you understand?

OpenStudy (anonymous):

yes can u answer another question

OpenStudy (anonymous):

What is the rate of change between the interval of x = pi over 4 and x = pi over two

OpenStudy (kropot72):

Good. Please post your other question as a new question, the reason being that I must log out now.

OpenStudy (anonymous):

ok